# Tight Algorithm for Connected Odd Cycle Transversal Parameterized by Clique-width

### 12 Feb 2024 v1

@@ 1 Introduction
Relying on the Cut&Count technique, tight bounds for connectivity problems were also derived for different parameters such as pathwidth [11], cutwidth [6], modular treewidth [23] and clique-width [21, 7]. Clique-width, as the concerned parameter in this paper, is defined by the smallest value $k$, such that a given graph can be built recursively from the disjoint union of $k$-labeled graphs, by allowing bicliques between vertices of different labels, and relabeling vertices. Clique-width is a more general parameter than treewidth, since cliques have clique-width at most $2$, and unbounded treewidth, while graphs of treewidth $k$ have bounded clique-width (by $2^{\mathcal{O}(k)}$) [9, 8].
@@
In Section 2 we provide preliminaries and introduce some notation. In Section 3 we define the notion of patterns, representation and pattern operations. In Section 4 we present the main result of this paper. We describe the dynamic programming algorithm, and prove its running time. In Section 5 we define action sequences, and use them to prove the correctness of the algorithm. Finally, in Section 6 we present our SETH based lower bound, proving the tightness of our algorithm. We conclude in Section 7 with open problems and final thoughts.
@@ 2 Preliminaries
  • Create vertex $i(v)$ for $i\in\mathbb{N}$. This operation constructs a graph containing a single vertex and assigns label $i$ to this vertex.
@@
Let $x\in\mathcal{V}$. The subtree rooted at $x$ induces a subexpression $\mu_{x}$. We define $G_{x}=G_{\mu_{x}}$, $V_{x}=V(G_{x})$, $E_{x}=E(G_{x})$, $\operatorname{lab}_{x}=\operatorname{lab}_{\mu_{x}}$, $\mathcal{T}_{x}=\mathcal{T}_{\mu_{x}}$, and $\mathcal{V}_{x}=\mathcal{V}_{\mu_{x}}$. We also define the set $\mathcal{V}^{C}_{x}$ as the set of all create nodes in $\mathcal{T}_{x}$, and $\mathcal{V}^{CJ}_{x}$ as the set of all create and join nodes in $\mathcal{T}_{x}$. Given a set $S\subseteq V$, we define $S_{x}=S\cap V_{x}$. Also for a mapping $g:S\rightarrow U$ to some set $U$, we denote by $g_{x}$ the mapping $g_{|S_{x}}$, the restriction of $g$ to vertices in $S\cap V_{x}$.
We call a clique-expression $\mu$ linear, if for each union node $x$ with children $x_{1}$ and $x_{2}$, i.e. $\mu_{x}=\mu_{x_{1}}\uplus\mu_{x_{2}}$, it holds that $\mu_{x_{2}}$ is a create operation. We say that $G$ is $k$-labeled, if it holds that $\operatorname{lab}(v)\leq k$ for all $v\in V$. We say that a clique-expression $\mu$ is a $k$-expression if $G_{x}$ is a $k$-labeled graph for all $x\in\mathcal{V}$. We define the (linear) clique-width of a graph $G$, denoted by $\operatorname{cw}(G)$ ($\operatorname{lcw}(G)$ for linear clique-width) as the smallest value $k$ such that there exists a (linear) $k$-expression $\mu$, with $G_{\mu}$ isomorphic to $G$. We can assume without loss of generality, that any given $k$-expression defining a graph $G=(V,E)$ uses at most $O(|V|)$ union operations, and at most $O(|V|k^{2})$ unary operations [2, 9].
@@
Along this work, let $G=(V,E)$ be the input graph and let $\overline{b}\in\mathbb{N}$ be the target value in the Connected Odd Cycle Transversal instance. Let $n=|V|$ and $m=|E|$. Let $\mu$ a $k$-expression of $G$ for some value $k\in\mathbb{N}$, and $\mathcal{T}$ be the corresponding syntax tree. Let $r$ be the root of $\mathcal{T}$.
@@ 3 Connectivity patterns
Then we define $\operatorname*{PREP}(p)=P_{r}$.
Clearly, each pattern of $\operatorname*{PREP}(p)$ results from $p$ by removing all non-singletons different from $Z_{p}$ from it, and adding some subset of $\operatorname{inc}(p)$ to $Z_{p}$. Hence, it holds that $|\operatorname*{PREP}(p)|\leq 2^{|\operatorname{inc}(p)|}$. Given a set $S\subseteq\mathscr{P}_{C}$, we define
@@ 4 The Algorithm
In this section we present the main results of this paper, an algorithm for the Connected Odd Cycle Transversal problem, a dynamic programming scheme over the node of $\mathcal{T}$. We start by defining the family of states $\mathscr{I}$, that we use to index the vectors computed by the recursive formulas of the dynamic programming algorithm.
@@ 4.1 Indices of the algorithm
The family $\mathscr{P}_{CS}$ is of a special interest in this paper, since we use it, similar to [7], to count connectivity patterns of partial solutions over the graphs $G_{x}$ for all nodes $x\in\mathcal{V}$. Note that a $CS$-pattern $p$ is uniquely defined by its labels set and its zero-set. Given $X,Y\subseteq U$ the labels set and the zero-set of $p$ respectively, with $Y\subseteq X$, it holds that
$p=\big{\{}\{u\}\colon u\in X\big{\}}\cup\big{\{}\{0\}\cup\{u\colon u\in Y\}\big{\}}.$
We define the ordering $\preceq_{CS}$ over $\mathscr{P}_{CS}$, where for two $CS$-patterns $p,q\in\mathscr{P}_{CS}$ it holds that $p\preceq_{CS}q$ if and only if $Z_{p}\subseteq Z_{q}$ and $\operatorname{label}(p)\subseteq\operatorname{label}(q)$. Clearly, $(\mathscr{P}_{CS},\preceq_{CS})$ defines a lattice, with $p\lor q=r$ where $\operatorname{label}(r)=\operatorname{label}(p)\cup\operatorname{label}(q)$, and $Z_{r}=Z_{p}\cup Z_{q}$. Hence, the join operation over this lattice corresponds to the union operation over $CS$-patterns $p\oplus q$.
@@
We fix two colors black $\boldsymbol{B}$ and white $\boldsymbol{W}$. Let $\overline{C}=\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W},\boldsymbol{BW}\}$ be the set of all unordered combinations of these two colors. We define the ordering $\sqsubseteq_{c}$ over $\overline{C}$ defined by inclusion over these combinations, namely, given by the two chains $\boldsymbol{\emptyset}\sqsubseteq_{c}\boldsymbol{B}\sqsubseteq_{c}\boldsymbol{BW}$ and $\boldsymbol{\emptyset}\sqsubseteq_{c}\boldsymbol{W}\sqsubseteq_{c}\boldsymbol{BW}$. We denote by $\sqcup$ the join operation, and by $\sqcap$ the meet operation over the underlying lattice. We call two colors $x,y\in\overline{C}$ consistent (denoted by $x\sim y$), if $x\sqcap y=\boldsymbol{\emptyset}$.
We denote the family of all colorings of $k$ labels by $\mathscr{C}=\overline{C}^{[k]}$. We define the ordering $\preceq_{C}$ over colorings, where for two colorings $c_{1},c_{2}\in\mathscr{C}$, it holds that $c_{1}\preceq_{C}c_{2}$ if it holds that $c_{1}(i)\sqsubseteq_{c}c_{2}(i)$ for all $i\in[k]$. Finally, we denote by $\oplus_{C}$ the join operation over the underlying lattice, i.e. for two colorings $c_{1},c_{2}$ and for all $i\in[k]$ it holds
@@
Indices of the dynamic programming algorithm
Let $\mathscr{I}=\mathscr{P}_{CS}\times\mathscr{C}$ be the family of all combinations of a $CS$-pattern and a coloring. We call a pair $(p,c)\in\mathscr{P}_{CS}\times\mathscr{C}$ a state, and $\mathscr{I}$ the family of all states. We index the dynamic programming tables by the elements of this family. While the pattern $p$ will indicate the connectivity of a solution, the coloring $c$ indicates the colors appearing in each label class in a witness of a partial solution, allowing to extend this partial solution correctly, keeping a proper 2-coloring of the rest of the graph.
@@ 4.2 Description of the Algorithm
  • Create node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c_{X}:[k]\rightarrow\overline{C}$ as
@@
    Let $p=[0i]$ if $v=v_{0}$, and $p=[i]$ otherwise. Let $p_{1}=\operatorname{forget}(p,i)$ and $p_{2}=\operatorname{fix}(p,i)$. We set all values $T{\big{[}{x,{b},{w}}\big{]}}\big{[}{q,c}\big{]}$ to zero, for all values of $q$ and $c$, and then we add one to the following indices:
@@
In the following, we show how to compute these tables in time $\mathcal{O}^{*}(12^{k})$. We assume that lookups and updates of a single element of any of these vectors can be done in logarithmic time in the number of indices and the size of the tables, and hence, in polynomial time in $n$.
Although computing the tables $T{\big{[}{x,{b},{w}}\big{]}}$ for a union node $x$ in the naive way requires polynomial time in ${(12^{2})}^{k}=144^{k}$, one can apply fast convolution methods to compute these tables more efficiently.
@@
We define the bijective mapping $\rho\colon\mathscr{I}\rightarrow\mathcal{L}^{*}$, as $\rho(p,c)=((x_{1},y_{1}),\dots(x_{k},y_{k}))$, where
$x_{i}=\begin{cases}3&\colon i\in Z_{p},\\ 2&\colon i\in\operatorname{label}(p)\setminus Z_{p},\\ 1&\colon\text{Otherwise,}\end{cases}$
and
$y_{i}=\begin{cases}1&\colon c(i)=\boldsymbol{\emptyset},\\ 2&\colon c(i)=\boldsymbol{B},\\ 3&\colon c(i)=\boldsymbol{W},\\ 4&\colon c(i)=\boldsymbol{BW}.\end{cases}$
@@
Let $(p_{1},c_{1}),(p_{2},c_{2})\in\mathscr{I}$. We show that $(p_{1},c_{1})\preceq(p_{2},c_{2})$ if and only if $\rho(p_{1},c_{1})\preceq\rho(p_{2},c_{2})$. Let $\rho(p_{1},c_{1})=(x^{1}_{1},y^{1}_{1},\dots,x^{1}_{k},y^{1}_{k})$ and $\rho(p_{2},c_{2})=(x^{2}_{1},y^{2}_{1},\dots,x^{2}_{k},y^{2}_{k})$. It holds that $p_{1}\preceq_{CS}p_{2}$ if and only if $Z_{p_{1}}\subseteq Z_{p_{2}}$ and $\operatorname{label}(p_{1})\subseteq\operatorname{label}(p_{2})$. This is the case if and only if for all $i\in[k]$ it holds that $x^{1}_{i}=3$ implies that $x^{2}_{i}=3$, and $x^{1}_{i}=2$ implies that $x^{2}_{i}\geq 2$, which is the case if and only if $x^{1}_{i}\geq x^{2}_{i}$.
On the other hand, it holds that $c_{1}\preceq_{C}c_{2}$ if and only if $c_{1}(i)\sqsubseteq_{c}c_{2}(i)$ for all $i\in[k]$, which is the case if and only if $c_{2}(i)=\boldsymbol{BW}$, $c_{1}(i)=\boldsymbol{\emptyset}$ or $c_{1}(i)=c_{2}(i)$, which is the case if and only if $y^{2}_{i}=\boldsymbol{BW}$, $y^{1}_{i}=\boldsymbol{\emptyset}$, or $y^{1}_{i}=y^{2}_{i}$ for all $i\in[k]$. Hence, $(p_{1},c_{1})\preceq(p_{2},c_{2})$ if and only if $(x^{1}_{i},y^{1}_{i})\preceq_{0}(x^{2}_{i},y^{2}_{i})$ for all $i\in[k]$, which is the case if and only if $\rho(p_{1},y_{1})\preceq\rho(p_{2},y_{2})$. ∎
@@
For a a fixed tuple $({b}_{1},{b}_{2},{w}_{1},{w}_{2})$, we define the tables $T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}\in\{0,1\}^{\mathscr{I}}$ by
$T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}\big{[}{(p,c)}\big{]}=\sum\limits_{\begin{subarray}{c}p_{1}\oplus p_{2}=p\\ c_{1}\oplus_{C}c_{2}=c\end{subarray}}T{\big{[}{x_{1},{b}_{1},{w}_{1}}\big{]}}\big{[}{(p_{1},c_{1})}\big{]}\cdot T{\big{[}{x_{2},{b}_{2},{w}_{2}}\big{]}}\big{[}{(p_{2},c_{2})}\big{]}.$
Clearly, the table $T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}$ is the $\lor$-product of $T{\big{[}{x_{1},{b}_{1},{w}_{1}}\big{]}}$ and $T{\big{[}{x_{2},{b}_{2},{w}_{2}}\big{]}}$ over $\mathscr{I}$. Hence, by Corollary 4.6, it can be computed in time $\mathcal{O}^{*}(12^{k})$. On the other hand, it holds that
@@
This clearly the case for a create node. For a relabel node, for all values of ${b}$ and ${w}$, we initialize $T{\big{[}{x,{b},{w}}\big{]}}$ to $\overline{0}$ and iterate over all states in $\mathscr{I}$. For each such pair $(p,c)$, we add $T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}$ to $T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p_{i\rightarrow j},c_{i\rightarrow j})}\big{]}$.
For a join node, for all values of ${b}$ and ${w}$ we initialize $T{\big{[}{x,{b},{w}}\big{]}}$ to $\overline{0}$. Then we iterate again over all values of ${b}$ and ${w}$, and over all pairs $(c,p)\in\mathscr{I}$. For each such pair, if $T\big{[}{x^{\prime},{b},{w}}\big{]}\big{[}{(p,c)}\big{]}=0$ or $c(i)$ and $c(j)$ are not consistent, we skip this pair. Otherwise, let $p^{\prime}=\square_{i,j}p$. For each $\ell\in[4]$, Let $p_{\ell}=\operatorname{Ac}(p^{\prime},\ell)$ and $P_{\ell}=\operatorname*{PREP}(p_{\ell})$. Then we add one to each $T\big{[}{x,{b},{w}+\mathbf{d}(x,i)}\big{]}\big{[}{(c,q)}\big{]}$ for each $q\in P_{\ell}$.
Now we show that $|P_{\ell}|\leq 3$ for all $\ell\in[4]$. If $\{i,j\}\not\subseteq\operatorname{label}(p^{\prime})$, then $p_{\ell}=p^{\prime}=p\in\mathscr{P}_{CS}$, and hence, $P_{\ell}=\{p^{\prime}\}$ for each $\ell\in[4]$. Otherwise, we have $\operatorname{inc}(p^{\prime})=\{i,j\}$. If $\{i,j\}\cap Z_{p}\neq\emptyset$, then it holds that $Z_{p^{\prime}}$ is the only non-singleton in $p^{\prime}$, and hence, $p_{\ell}\in\mathscr{P}_{CS}$ for all $\ell\in[4]$. Otherwise, it holds that $\{i,j\}$ is the only non-singleton different from $Z_{p^{\prime}}$ in $p^{\prime}$. Hence, it holds that $|P_{1}|=3$ and $p_{\ell}\in\mathscr{P}_{CS}$ for $\ell\in\{2,3,4\}$.

### 26 Feb 2024 v2

@@ 1 Introduction
Relying on the Cut&Count technique, tight bounds for connectivity problems were also derived for different parameters such as pathwidth [11], cutwidth [6], modular treewidth [23] and clique-width [21, 7]. In this paper, we focus on Clique-width, a graph parameter defined by the smallest value $k$, such that a given graph can be built recursively from the disjoint union of $k$-labeled graphs, by allowing bicliques between vertices of different labels, and relabeling vertices. Clique-width is a more general parameter than treewidth, since cliques have clique-width at most $2$, and unbounded treewidth, while graphs of treewidth $k$ have bounded clique-width (at most $2^{\mathcal{O}(k)}$) [9, 8].
@@
In Section 2 we provide preliminaries and introduce some notation. In Section 3 we define the notion of patterns, representation and pattern operations. In Section 4 we present the main result of this paper. We describe the dynamic programming algorithm, and prove its running time. In Section 5 we define action-sequences, and use them to prove the correctness of the algorithm. Finally, in Section 6 we present our SETH based lower bound, proving the tightness of our algorithm. We conclude in Section 7 with open problems and final thoughts.
@@ 2 Preliminaries
  • Introduce vertex $i(v)$ for $i\in\mathbb{N}$. This operation constructs a graph containing a single vertex and assigns label $i$ to this vertex.
@@
Let $x\in\mathcal{V}$. The subtree rooted at $x$ induces a subexpression $\mu_{x}$. We define $G_{x}=G_{\mu_{x}}$, $V_{x}=V(G_{x})$, $E_{x}=E(G_{x})$, $\operatorname{lab}_{x}=\operatorname{lab}_{\mu_{x}}$, $\mathcal{T}_{x}=\mathcal{T}_{\mu_{x}}$, and $\mathcal{V}_{x}=\mathcal{V}_{\mu_{x}}$. We also define the set $\mathcal{V}^{C}_{x}$ as the set of all introduce nodes in $\mathcal{T}_{x}$, and $\mathcal{V}^{CJ}_{x}$ as the set of all introduce and join nodes in $\mathcal{T}_{x}$. Given a set $S\subseteq V$, we define $S_{x}=S\cap V_{x}$. Also for a mapping $g:S\rightarrow U$ to some set $U$, we denote by $g_{x}$ the mapping $g_{|S_{x}}$, the restriction of $g$ to vertices in $S\cap V_{x}$.
We call a clique-expression $\mu$ linear, if for each union node $x$ with children $x_{1}$ and $x_{2}$, i.e. $\mu_{x}=\mu_{x_{1}}\uplus\mu_{x_{2}}$, it holds that $\mu_{x_{2}}$ is an introduce operation. We say that $G$ is $k$-labeled, if it holds that $\operatorname{lab}(v)\leq k$ for all $v\in V$. We say that a clique-expression $\mu$ is a $k$-expression if $G_{x}$ is a $k$-labeled graph for all $x\in\mathcal{V}$. We define the (linear) clique-width of a graph $G$, denoted by $\operatorname{cw}(G)$ ($\operatorname{lcw}(G)$ for linear clique-width) as the smallest value $k$ such that there exists a (linear) $k$-expression $\mu$, with $G_{\mu}$ isomorphic to $G$. We can assume without loss of generality, that any given $k$-expression defining a graph $G=(V,E)$ uses at most $O(|V|)$ union operations, and at most $O(|V|k^{2})$ unary operations [2, 9].
@@
Clearly, $\Delta$ is an associative operation. Hence, for a family of sets =$\mathcal{F}=\{S_{1},\dots,S_{n}\}$, the following notion is well-defined
$\limits_{S\in\mathcal{F}}S=S_{1}\Delta S_{2}\Delta\dots\Delta S_{n}.$
@@
Conventions.
Along this work, let $G=(V,E)$ be the input graph and let $\overline{b}\in\mathbb{N}$ be the target value in the Connected Odd Cycle Transversal instance. Let $n=|V|$ and $m=|E|$. Let $\mu$ a $k$-expression of $G$ for some value $k\in\mathbb{N}$, and $\mathcal{T}$ be the corresponding syntax tree. Let $r$ be the root of $\mathcal{T}$. A partial solution at $x\in\mathcal{V}$ is a set of vertices $S\subseteq V_{x}$.
For a function $f$, we denote by $\mathcal{O}^{*}(f(k))$ the running time $\mathcal{O}(f(k)\cdot\operatorname*{Poly}(n)),$ where $\operatorname*{Poly}(n)$ denotes some arbitrary polynomial in $n$. When we say that some function is bounded polynomially, we always mean polynomially in $n$ (and hence, in the size of the input).
@@ 3 Connectivity patterns
Then we define $\operatorname*{PREP}(p)=P_{r}$. Given a set $S\subseteq\mathscr{P}_{C}$, we define
@@
Observation 3.11.
For a complete pattern $p\in\mathscr{P}_{C}$, it holds that each pattern of $\operatorname*{PREP}(p)$ results from $p$ by removing all non-singleton sets different from $Z_{p}$, and adding some subset of $\operatorname{inc}(p)$ to $Z_{p}$. Hence, it holds that $|\operatorname*{PREP}(p)|\leq 2^{|\operatorname{inc}(p)|}$.
@@ 4 The Algorithm
In this section we present the main result of this paper, namely a dynamic programming algorithm over $\mathcal{T}$ solving the Connected Odd Cycle Transversal problem. We start by defining the family of states $\mathscr{I}$ that we use to index the vectors computed by the recursive formulas of the dynamic programming algorithm.
@@ 4.1 Indices of the algorithm
The family $\mathscr{P}_{CS}$ is of a special interest in this paper, since we use it, similar to [7], to count connectivity patterns of partial solutions over the graphs $G_{x}$ for all nodes $x\in\mathcal{V}$. Note that a $CS$-pattern $p$ is uniquely defined by its set of labels and its zero-set; Given $X,Y\subseteq U$ the set of labels and the zero-set of $p$ respectively, where $Y\subseteq X$, it holds that
$p=\big{\{}\{u\}\colon u\in X\big{\}}\cup\big{\{}\{0\}\cup Y\big{\}}.$
We define the partial ordering $\preceq_{CS}$ over $\mathscr{P}_{CS}$, where for two $CS$-patterns $p,q\in\mathscr{P}_{CS}$ it holds that $p\preceq_{CS}q$ if and only if $Z_{p}\subseteq Z_{q}$ and $\operatorname{label}(p)\subseteq\operatorname{label}(q)$. Clearly, $(\mathscr{P}_{CS},\preceq_{CS})$ defines a lattice, with $p\lor q=r$ where $\operatorname{label}(r)=\operatorname{label}(p)\cup\operatorname{label}(q)$, and $Z_{r}=Z_{p}\cup Z_{q}$. Hence, the join operation over this lattice corresponds to the union operation over $CS$-patterns $p\oplus q$.
@@
We define the set $C_{0}=\{\boldsymbol{B},\boldsymbol{W}\}$, and call it the set of basic colors, where $\boldsymbol{B}$ stands for black, and $\boldsymbol{W}$ stands for white. We define the set of colors $\overline{C}=\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W},\boldsymbol{BW}\}$, where each element of $\overline{C}$ corresponds to a different subset of $C_{0}$ in the natural way. We define the ordering $\sqsubseteq_{c}$ over $\overline{C}$ given by inclusion over the corresponding subsets, namely, given by the two chains $\boldsymbol{\emptyset}\sqsubseteq_{c}\boldsymbol{B}\sqsubseteq_{c}\boldsymbol{BW}$ and $\boldsymbol{\emptyset}\sqsubseteq_{c}\boldsymbol{W}\sqsubseteq_{c}\boldsymbol{BW}$. We denote by $\sqcup$ the join operation, and by $\sqcap$ the meet operation over the underlying lattice. We call two colors $x,y\in\overline{C}$ consistent (denoted by $x\sim y$), if $x\sqcap y=\boldsymbol{\emptyset}$.
We call a mapping $c:U\rightarrow\overline{C}$ a coloring. We denote the family of all colorings of $U$ by $\mathscr{C}=\overline{C}^{U}$. We define the ordering $\preceq_{C}$ over colorings, where for two colorings $c_{1},c_{2}\in\mathscr{C}$, it holds that $c_{1}\preceq_{C}c_{2}$ if it holds that $c_{1}(i)\sqsubseteq_{c}c_{2}(i)$ for all $i\in U$. Finally, we denote by $\oplus_{C}$ the join operation over the underlying lattice, i.e. for $c_{1},c_{2}\in\mathscr{C}$ and for all $i\in U$ it holds
@@
Indices of the dynamic programming routine
Let $\mathscr{I}=\mathscr{P}_{CS}\times\mathscr{C}$ be the family of all pairs of a $CS$-pattern and a coloring. We call a pair $(p,c)\in\mathscr{P}_{CS}\times\mathscr{C}$ a state, and $\mathscr{I}$ the family of all states. We index the dynamic programming tables by the elements of this family. While the pattern $p$ indicates the connectivity of a partial solution, the coloring $c$ indicates the basic colors appearing in each label class in a witness of this partial solution, allowing to extend this partial solution correctly, preserving a proper 2-coloring of the rest of the graph.
@@ 4.2 Description of the Algorithm
Intuitively, actions (Definition 3.8) allow to build representations (in an existential sense) of different weights of all partial solutions by a dynamic programming scheme over $\mathcal{T}$ using complete patterns only. However, since the number of complete patterns is at least the number of all partitions of all subsets of $U$, we seek to reduce the space of the states of the dynamic programming routine, by representing (in a count-preserving manner) these patterns using $CS$-patterns only (Definition 3.10).
In this section, we define the vectors $T{\big{[}{x,{b},{w}}\big{]}}\in\{0,1\}^{\mathscr{I}}$ for $x\in\mathcal{V}$ and all values ${b}$ and ${w}$. These vectors constitute the dynamic programming tables of our algorithm. We show that these vectors can be computed in time $\mathcal{O}^{*}(12^{k})$. In the following sections, we will prove the correctness of the algorithm in two stages. In the first stage we show that any pattern, corresponding to a partial solution at a node $x$, can be represented by a set of complete patterns that can be built using different actions at different nodes in $\mathcal{V}^{CJ}_{x}$. We assign different weights to different actions at each node of $\mathcal{T}$. In the second stage, we show that the tables $T{\big{[}{x,{b},{w}}\big{]}}$ count for each $CS$-pattern p, the number of representations of weight $w$ of solutions of size ${b}$ that are consistent with $p$. By appropriately choosing the weight function, we show that this suffices to solve the Connected Odd Cycle Transversal problem with hight probability.
@@
  • Introduce node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c_{X}:[k]\rightarrow\overline{C}$ as
@@
    Let $p=[0i]$ if $v=v_{0}$, and $p=[i]$ otherwise. Let $p_{1}=\operatorname{forget}(p,i)$ and $p_{2}=\operatorname{fix}(p,i)$. We set all values $T{\big{[}{x,{b},{w}}\big{]}}\big{[}{q,c}\big{]}$ to zero, for all values of $q$ and $c$, and then we add one to each of the following entries:
@@
In the following, we show how to compute these tables in time $\mathcal{O}^{*}(12^{k})$. We assume that lookups and updates of a single entry of any of these vectors can be done in time logarithmic in the number of indices, and hence, in polynomial time in $n$.
Although computing the tables $T{\big{[}{x,{b},{w}}\big{]}}$ for a union node $x$ in the naive way requires time polynomial in ${(12^{k})}^{2}=144^{k}$, we show that one can apply fast convolution methods to compute these tables more efficiently.
@@
We define the bijective mapping $\rho\colon\mathscr{I}\rightarrow\mathcal{L}^{*}$, as $\rho(p,c)=((x_{1},y_{1}),\dots(x_{k},y_{k}))$, where for $i\in[k]$ it holds
[…]
@@
Let $(p_{1},c_{1}),(p_{2},c_{2})\in\mathscr{I}$. We show that $(p_{1},c_{1})\preceq(p_{2},c_{2})$ if and only if $\rho(p_{1},c_{1})\preceq\rho(p_{2},c_{2})$. Let $\rho(p_{1},c_{1})=(x^{1}_{1},y^{1}_{1},\dots,x^{1}_{k},y^{1}_{k})$ and $\rho(p_{2},c_{2})=(x^{2}_{1},y^{2}_{1},\dots,x^{2}_{k},y^{2}_{k})$. It holds that $p_{1}\preceq_{CS}p_{2}$ if and only if $Z_{p_{1}}\subseteq Z_{p_{2}}$ and $\operatorname{label}(p_{1})\subseteq\operatorname{label}(p_{2})$. This is the case if and only if for all $i\in[k]$ it holds that $x^{1}_{i}=3$ implies that $x^{2}_{i}=3$, and $x^{1}_{i}=2$ implies that $x^{2}_{i}\geq 2$, which holds if and only if $x^{1}_{i}\leq x^{2}_{i}$.
On the other hand, it holds that $c_{1}\preceq_{C}c_{2}$ if and only if $c_{1}(i)\sqsubseteq_{c}c_{2}(i)$ for all $i\in[k]$, which is the case if and only if $c_{2}(i)=\boldsymbol{BW}$, $c_{1}(i)=\boldsymbol{\emptyset}$ or $c_{1}(i)=c_{2}(i)$, which is the case if and only if $y^{2}_{i}=4$, $y^{1}_{i}=1$, or $y^{1}_{i}=y^{2}_{i}$ for all $i\in[k]$. Hence, $(p_{1},c_{1})\preceq(p_{2},c_{2})$ if and only if $(x^{1}_{i},y^{1}_{i})\preceq_{0}(x^{2}_{i},y^{2}_{i})$ for all $i\in[k]$, which is the case if and only if $\rho(p_{1},y_{1})\preceq\rho(p_{2},y_{2})$. ∎
@@
For a fixed tuple $({b}_{1},{b}_{2},{w}_{1},{w}_{2})$, we define the tables $T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}\in\{0,1\}^{\mathscr{I}}$ by
$T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}\big{[}{(p,c)}\big{]}=\sum\limits_{\begin{subarray}{c}p_{1}\oplus p_{2}=p\\ c_{1}\oplus_{C}c_{2}=c\end{subarray}}T{\big{[}{x_{1},{b}_{1},{w}_{1}}\big{]}}\big{[}{(p_{1},c_{1})}\big{]}\cdot T{\big{[}{x_{2},{b}_{2},{w}_{2}}\big{]}}\big{[}{(p_{2},c_{2})}\big{]},$
i.e., the table $T\big{[}{x,({b}_{1},{b}_{2},{w}_{1},{w}_{2})}\big{]}$ is the $\lor$-product of $T{\big{[}{x_{1},{b}_{1},{w}_{1}}\big{]}}$ and $T{\big{[}{x_{2},{b}_{2},{w}_{2}}\big{]}}$ over $\mathscr{I}$. Hence, by Corollary 4.6, it can be computed in time $\mathcal{O}^{*}(12^{k})$. On the other hand, it holds that
@@
This is clearly the case for an introduce node. For a relabel node, for all values of ${b}$ and ${w}$, we initialize $T{\big{[}{x,{b},{w}}\big{]}}$ to $\overline{0}$ and iterate over all states in $\mathscr{I}$. For each such pair $(p,c)$, we add $T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}$ to $T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p_{i\rightarrow j},c_{i\rightarrow j})}\big{]}$.
For a join node, for all values of ${b}$ and ${w}$ we initialize $T{\big{[}{x,{b},{w}}\big{]}}$ to $\overline{0}$. Then we iterate again over all values of ${b}$ and ${w}$, and over all pairs $(p,c)\in\mathscr{I}$. For each such pair, if $T\big{[}{x^{\prime},{b},{w}}\big{]}\big{[}{(p,c)}\big{]}=0$ or $c(i)$ and $c(j)$ are not consistent, we skip this pair. Otherwise, let $p^{\prime}=\square_{i,j}p$. For each $\ell\in[4]$, let $p_{\ell}=\operatorname{Ac}(p^{\prime},\ell)$ and $P_{\ell}=\operatorname*{PREP}(p_{\ell})$. Then we add one to each entry $T\big{[}{x,{b},{w}+\mathbf{d}(x,\ell)}\big{]}\big{[}{(q,c)}\big{]}$ for each $q\in P_{\ell}$.
Since $\{i,j\}$ is the only non-singleton set different from $Z_{p_{\ell}}$ that can appear in $p_{\ell}$, it holds that $\operatorname{inc}(p_{\ell})\subseteq\{i,j\}$. Hence, it holds by 3.11 that $|P_{\ell}|\leq 4$. In total, the algorithm performs at most 16 binary addition operations for each fixed values of ${b}$ and ${w}$, and for each pair $(p,c)\in\mathscr{I}$.